OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Sn: 2 is a factor of n2 + 7n

OpenStudy (anonymous):

\[Sn=n ^{2}+7n\] \[S1=1^{2}+7\times1=1+7=8=2\times4,\] which shows S1 is true. \[S2=2^{2}+7\times2=4+14=18=2\times9\] which shows S2 is true. \[S3=3^{2}+7\times3=9+21=30=2\times15\] which shows S3 is true. Here S1,S2,S3 are all divisible by 2 ,hence all are true.

OpenStudy (anonymous):

You are the most awesome. Thanks!